# Hello Again! LLM-powered Personalized Agent for Long-term Dialogue

### v1

@@ 1 Introduction
Open-domain dialogue systems aim to establish long-term, personalized interactions with users via human-like chatbots [1, 2, 3]. Unlike most existing studies [4, 5, 6] that are limited to brief, single-session interactions spanning 2-15 turns, real-life scenarios often necessitate a chatbot’s capability for long-term companionship and familiarity [1, 2, 7]. Achieving this requires the chatbot not only to understand and remember extensive dialogue histories but also to faithfully reflect and consistently update both the user’s and its personalized characteristics [1, 7, 8].
Motivated by real-life demands, the core challenge of open-domain dialogue systems is to simultaneously maintain long-term event memory and preserve persona consistency [9, 10, 11, 3]. Existing research often addresses these aspects separately—focusing either on event memory or persona extraction—thereby hindering long-term consistency. Current strategies for event memory typically involve constructing a memory bank that stores historical event summaries, complemented by retrieval-augmented approaches to access relevant information for response generation [12, 13]. Studies on persona-based dialogue rang from unidirectional user modeling [14] to bidirectional agent-user modeling [15, 16, 3], enhancing personalized chat abilities by leveraging profile information. Worse still, the aforementioned methods are highly dependent on specific model architectures, making them challenging to adapt to other models. Additionally, These dialogue models largely lack zero-shot generalization capabilities, essential for effective deployment across various real-world domains [2, 3]. We conjecture that an optimal long-term dialogue framework should be model-agnostic, deployable in various real-world domains, and capable of autonomously integrating comprehensive data from both event memories and personas, as illustrated in Figure 1. However, developing such a model-agnostic, cross-domain, and autonomous framework remains unexplored and challenging.
@@
Benefiting from the excellent human-like cognitive and reasoning abilities of large language models (LLM), there is an increasing trend [17, 18, 19, 20, 21] to employ LLMs as the cores of agent-based simulation systems to automate the process of perception, decision-making, and problem-solving. While recent studies have developed LLM-powered agents in various fields, such as economics [22], politics [23], sociology [24], and recommendation [21], its application in open-domain dialogue remains unexplored. To effectively support long-term open-domain dialogue, an LLM-powered dialogue agent framework should exhibit broad generality, cross-domain adaptability, and the ability to dynamically refine information across dimensions like events, user personalities, and agent personalities.
@@
We conduct comprehensive experiments on two illustrative long-term multi-session daily dialogue datasets, MSC [1] and Conversation Chronicles (CC) [7], to evaluate the effectiveness, generality, and cross-domain capabilities of the proposed framework. In terms of effectiveness, LD-Agent achieves state-of-the-art performance on both benchmarks, significantly outperforming existing methods [2, 25, 26]. To assess generality, we examine the framework from both model and task perspectives. From the model perspective, LD-Agent is evaluated across a range of both online and offline models, including LLMs [25] and non-LLMs [26]. From the task perspective, we extend our evaluation to multiparty dialogue tasks [27], where LD-Agent also demonstrates substantial improvements, showcasing its adaptability across different models and tasks. Regarding the method’s cross-domain capabilities, we design two cross-domain settings: tuning the model on the MSC dataset and testing it on the CC dataset, and vice versa. In both scenarios, LD-Agent shows competitive performance, nearly matching the results of in-domain training.
@@
  • We develop LD-Agent, a general long-term dialogue agent framework, considering both historical events and personas. The event memory module ensures dialogue coherence across sessions, while the persona module ensures character consistency.
@@ 2 Method
In this section, we introduce the detailed description of LD-Agent with the framework shown in Figure 2. We first introduce the task definition of long-term dialogue in Section. 2.1. Consequently, we separately introduce the mechanism of event perception (Section. 2.2), dynamic personas extraction (Section. 2.3), and response generation (Section. 2.4).
@@ 2.2 Event Perception
The event memory module is designed to perceive historical events to generate coherent responses across interval time. As shown in Figure 2, this event memory module is segmented into two major sub-modules that focus separately on long-term and short-term memory.
@@ 2.2.1 Long-term Memory
The long-term memory module aims to extract and encode events from past sessions. Specifically, this involves recording the occurrence times $t$ and brief summaries $o$ into representations that are stored in a low-cost memory bank $M_{L}=\{\phi(t_{j},o_{j})\mid j\in\{1,2,\dots,l\}\}$. Here, $\phi(\cdot)$ indicates the text encoder (e.g., MiniLM [28]), and $l$ specifies the length of the memory bank. The encoded representations are then efficiently retrieved through an embedding-based mechanism, which enhances the accessibility of the stored memory.
@@
Different from previous agent approaches [20, 21, 29] that entirely rely on LLM’s zero-shot ability to excavate and summarize events, we apply instruction tuning [30] to the event summary module, which can directly improve the event summary quality. Specifically, we rebuild the DialogSum dataset [31], a large-scale dialogue summarization dataset, into the following format: (1) an introduction to the task background, (2) the related conversations that need to be understood, and (3) detailed summarization requests. These three parts serve as input prompts (see Appendix. D.1 for more details), combined with the original summaries from DialogSum as answers, and are jointly used to fine-tune the event summary module, thereby directly improving the quality of event summarization.
@@
To improve retrieval accuracy, we employ a retrieval mechanism that comprehensively considers semantic relevance, topic overlap, and time decay. Optimizing the retrieval accuracy of agent memory is challenging due to the difficulty in obtaining accurate memory retrieval data. Most existing methods [20, 21] use event summaries as keys and context as queries, calculating the query-key semantic relevance score $s_{\text{sem}}$ to find relevant memories, which inevitably results in significant errors. To enhance retrieval reliability, we extract nouns from corresponding conversations with the summaries to construct a topic library $V$ and calculate topic overlap score $s_{\text{top}}$ by:
$s_{\text{top}}=\frac{1}{2}(\frac{\left|V_{q}\cap V_{k}\right|}{V_{q}}+\frac{\left|V_{q}\cap V_{k}\right|}{V_{k}}),$ (1)
where $V_{q}$, $V_{k}$ denote the topic noun set of query and key. Additionally, we apply a time decay coefficient $\lambda_{t}=e^{{-t}/{\tau}}$ to reweight the overall retrieval score $s_{r}$, signified as:
@@
To avoid retrieving inappropriate memory due to no suitable memories existing, we implement a semantic threshold $\gamma$. Only memories with semantic score $s_{\text{sem}}$ greater than $\gamma$ could be retrieved. If no appropriate memories are retrieved, “No relevant memory” will be returned. Eventually, the process of retrieving relevant memory $m$ can be denoted as:
$m=\psi(M_{L},\gamma).$ (3)
@@ 2.2.2 Short-term Memory
The short-term memory module actively manages a dynamic dialogue cache $M_{S}=\{(t_{i},u_{i})|i=\{1,2,3,\dots,r_{c}\}\}$ with timestamps to preserve the detailed context of the current session. Upon receiving a new utterance $u^{\prime}$, the module first evaluates the time interval between the current time $t^{\prime}$ and the last recorded time $t_{r_{c}}$ in the cache. If this interval exceeds a threshold $\beta$, the module triggers the long-term memory module to summarize the cached dialogue entries, creating new event records for storage in the long-term memory bank. Simultaneously, the short-term memory cache is cleared, and the new dialogue record $(t^{\prime},u^{\prime})$ is added to the cache. The mathematical expression of this process is given by:
@@
where $M^{\prime}_{L}$ denotes the updated long-term memory bank, $o=A(\cdot)$ represents the event summary function, which process the accumulated dialogue in $M_{S}$.
@@ 2.3 Dynamic Personas Extraction
The persona module is pivotal in maintaining long-term persona consistency for both participants in a dialogue system. Drawing inspiration from prior work [3], we adopt a bidirectional user-agent modeling approach, utilizing a tunable persona extractor to manage long-term persona bank $P_{u}$ and $P_{a}$ for the user and agent, respectively. Specifically, we develop an open-domain, utterance-based persona extraction dataset derived from MSC [1]. We enhance the persona extractor with LoRA-based instruction tuning, which allows for the dynamic extraction of personality traits during conversations. These traits are subsequently stored in the corresponding character’s persona bank. For utterances devoid of personality traits, the module outputs “No Trait”. Additionally, we employ a tuning-free strategy that harnesses the zero-shot capabilities of LLM models to directly extract personas based on prompts. To further improve the agent’s ability to excavate user personas without training, we adjust our reasoning strategy from direct reasoning to a Chain-of-Thought reasoning [32] (see Appendix. D.2).
@@ 2.4 Response Generation
Upon receiving a new user utterance $u^{\prime}$, the agent integrates various inputs: retrieved relevant memories $m$, short-term context $M_{S}$, and the personas $P_{u}$ and $P_{a}$ for the user and agent, respectively. These combined inputs are fed into a response generator to deduce an appropriate response $r$, as formulated in Eq. 5.
@@
To enhance the agent’s ability for coherent and contextually appropriate responses, we develop a long-term, multi-session dialogue dataset, featuring dynamic retrieval memories, context, and personas sourced from the MSC and CC datasets for generator tuning. Specifically, for each sample, covering five sessions, we dynamically simulate the entire progression of the conversation. As each new utterance is introduced, the previously tuned modules for event summarization, persona extraction, and topic-aware memory retrieval are utilized to collect the necessary context, retrieved memories, and both user and agent personas related to the utterance. This comprehensive data is then integrated into a response generation prompt (see Appendix. D.3). The original responses from the MSC and CC datasets are used as ground truth sentences.
@@ 3.1 Evaluation Settings
In this subsection, we briefly introduce the experimental dataset, evaluation metrics, and baseline models in our study. Detailed evaluation settings are elaborated in Appendix. C.
Datasets.
Extensive experiments are conducted on two illustrative multi-session datasets, MSC [1] and CC [7], each comprising 5 sessions with approximately 50 conversational turns per sample, to investigate the effectiveness of LD-Agent on long-term dialogue scenarios. The experiments cover model independence assessment, module ablation, persona extractor analysis, and cross-domain evaluation.
Additionally, to evaluate the transferability of the LD-Agent, we apply our method to the Ubuntu IRC benchmark [27], a dataset known for its multiparty interaction tasks.
Metrics.
Our evaluation combines both automatic and human assessments to thoroughly investigate the effectiveness of LD-Agent. For automatic evaluation, we use three widely used standard metrics: BLEU-N (BL-N) [33], ROUGE-L (R-L) [34], and METEOR (MET) [35] to measure the quality of response generation. Additionally, accuracy (ACC) is employed to evaluate the classification performance of the persona extractor. In human evaluation, we measure topic coherence across sessions, interaction fluency, and user engagement using the metrics of coherence, fluency, and engagingness, respectively.
Baselines.
To demonstrate the effectiveness and model independence of LD-Agent, we deploy LD-Agent on multiple platforms and models. Specifically, the LLM-based models (online model: ChatGPT; offline model: ChatGLM [25]) and traditional language models (BlenderBot [26], and BART [36]) are employed as our baselines. In our experiments, The notation “ModelLDA” denotes models that incorporate the LD-Agent framework, while “Model” refers to the original baseline models without LD-Agent. Additionally, we also utilize HAHT [2], the previous state-of-the-art model in long-term dialogue task, as a contrast. See the above baselines stand and their role in rich literature in Appendix. A.
Table 1: Experimental results of the automatic evaluation for response generation on MSC and CC.
@@
Tuning ChatGLMLDA 10.70 5.63 23.31 10.03 5.12 21.55 9.07 4.06 20.19 8.96 4.01 19.94

### v2

@@ Abstract
Hello Again! LLM-powered Personalized Agent for Long-term Dialogue
Hao Li1thanks: These authors contribute equally to this work., Chenghao Yang2, An Zhang1thanks: An Zhang is the corresponding author., Yang Deng3, Xiang Wang2, Tat-Seng Chua1 1National University of Singapore 2University of Science and Technology of China 3Singapore Management University 18th.leolee@gmail.com, yangchenghao@mail.ustc.edu.cn, anzhang@u.nus.edu, ydeng@smu.edu.sg, xiangwang123@gmail.com, dcscts@nus.edu.sg
@@ 1 Introduction
Open-domain dialogue systems aim to establish long-term, personalized interactions with users via human-like chatbots Xu et al. (2022a); Zhang et al. (2022); Xu et al. (2022b). Unlike most existing studies Li et al. (2017); Zhang et al. (2018); Rashkin et al. (2019) that are limited to brief, single-session interactions spanning 2-15 turns, real-life scenarios often necessitate a chatbot’s capability for long-term companionship and familiarity Xu et al. (2022a); Zhang et al. (2022); Jang et al. (2023). Achieving this requires the chatbot not only to understand and remember extensive dialogue histories but also to faithfully reflect and consistently update both the user’s and its personalized characteristics.
Motivated by real-life demands, the core challenge of open-domain dialogue systems is to simultaneously maintain long-term event memory and preserve persona consistency Gu et al. (2019); Cao et al. (2022); Zhao et al. (2023); Xu et al. (2022b). Existing research often addresses these aspects separately—focusing either on event memory or persona extraction—thereby hindering long-term consistency. Current strategies for event memory typically involve constructing a memory bank that stores historical event summaries, complemented by retrieval-augmented approaches to access relevant information for response generation Chen et al. (2019); Zhang et al. (2019). Studies on persona-based dialogue rang from unidirectional user modeling Chen et al. (2023a) to bidirectional agent-user modeling Wu et al. (2020a); Liu et al. (2020); Xu et al. (2022b), enhancing personalized chat abilities by leveraging profile information. Worse still, the aforementioned methods are highly dependent on specific model architectures, making them challenging to adapt to other models. Additionally, These dialogue models largely lack zero-shot generalization capabilities, essential for effective deployment across various real-world domains Zhang et al. (2022); Xu et al. (2022b). We conjecture that an optimal long-term dialogue framework should be model-agnostic, deployable in various real-world domains, and capable of autonomously integrating comprehensive data from both event memories and personas, as illustrated in Figure 1. However, developing such a model-agnostic, cross-domain, and autonomous framework remains unexplored and challenging.
@@
Benefiting from the excellent human-like cognitive and reasoning abilities of large language models (LLM), there is an increasing trend Deng et al. (2023); Wang et al. (2023a); Qian et al. (2023); Park et al. (2023); Zhang et al. (2023a) to employ LLMs as the cores of agent-based simulation systems to automate the process of perception, decision-making, and problem-solving. While recent studies have developed LLM-powered agents in various fields, such as economics Cheng and Chin (2024), politics Hua et al. (2023), sociology Xu et al. (2024), and recommendation Zhang et al. (2023a), its application in open-domain dialogue remains unexplored. To effectively support long-term open-domain dialogue, an LLM-powered dialogue agent framework should exhibit broad generality, cross-domain adaptability, and the ability to dynamically refine information across dimensions like events, user personalities, and agent personalities.
@@
We conduct comprehensive experiments on two illustrative long-term multi-session daily dialogue datasets, MSC Xu et al. (2022a) and Conversation Chronicles (CC) Jang et al. (2023), to evaluate the effectiveness, generality, and cross-domain capabilities of the proposed framework. In terms of effectiveness, LD-Agent achieves state-of-the-art performance on both benchmarks, significantly outperforming existing methods Zhang et al. (2022); Zeng et al. (2023); Roller et al. (2021). To assess generality, we examine the framework from both model and task perspectives. From the model perspective, LD-Agent is evaluated across a range of both online and offline models, including LLMs Zeng et al. (2023) and non-LLMs Roller et al. (2021). From the task perspective, we extend our evaluation to multiparty dialogue tasks Hu et al. (2019), where LD-Agent also demonstrates substantial improvements, showcasing its adaptability across different models and tasks. Regarding the method’s cross-domain capabilities, we design two cross-domain settings: tuning the model on the MSC dataset and testing it on the CC dataset, and vice versa. In both scenarios, LD-Agent shows competitive performance, nearly matching the results of in-domain training.
@@
  • We develop LD-Agent, a general long-term dialogue agent framework, considering both historical events, ensuring coherence across sessions and personas, ensuring character consistency.
@@
2 Related Work
Long-term Dialogues.
Open-domain dialogue aims to develop a human-like chatbot that can emulate human conversation, facilitating free-flowing dialogue on a wide range of topics. However, the dialogue’s extent in earlier studies is often limited by conversation length, focusing primarily on brief conversations of about 2-15 turns within a single session Li et al. (2017); Zhang et al. (2018); Rashkin et al. (2019). To support more realistic and extended conversations, a series of studies have explored the role of both external Wang et al. (2023b, 2024) and internal knowledge Zhang et al. (2022); Xu et al. (2022b) on maintaining the feasibility of long-term dialogue. Commonly referenced external knowledge, such as commonsense Wang et al. (2024), medical Chen et al. (2023b), and psychological Chen et al. (2023c) knowledge, serves as supplementary guidance for the reasoning process, ensuring logical coherence in extended contexts. In parallel, internal knowledge captured dynamically during long conversations generally contains historical events Xu et al. (2022a); Zhang et al. (2023b, 2022); Jang et al. (2023) and personas Gu et al. (2019); Xu et al. (2022b); Cao et al. (2022); Deng et al. (2022). Historical events are typically summarized and stored into a memory bank to maintain dialogue coherence across sessions, while interlocutors’ personas are maintained via a dynamic persona memory bank, ensuring character consistency in long-term conversations. In this study, we focus on the internal knowledge to integrate dynamically updated historical events and personas to conduct long-term personalized conversations.
LLM-based Autonomous Agents.
AI Agent conception is geared towards autonomous environmental perception, decision-making, and problem-solving capabilities. With the large language models (LLMs) underlining their impressive generalization potential, leading to their widespread adoption as substitutes for human operators in various research fields Deng et al. (2023); Qian et al. (2023); Dillion et al. (2023); Zhang et al. (2023a); Park et al. (2023). Generally, these agents can be categorized into task-oriented agents Deng et al. (2023); Wang et al. (2023a); Qian et al. (2023); Zhong et al. (2024) and simulation-oriented agents Dillion et al. (2023); Shaikh et al. (2023); Gao et al. (2023a); Zhang et al. (2023a); Huang et al. (2023). Task-oriented agents are designed to accurately perform predefined tasks, as seen in applications for web assistance Deng et al. (2023), game-playing Wang et al. (2023a), and software development Qian et al. (2023). Conversely, simulation-oriented agents are devised to emulate human emotive and cognitive behaviors, having played roles in psychological studies Dillion et al. (2023), social networking platforms Gao et al. (2023a), conflict resolution scenarios Shaikh et al. (2023), and recommendation systems Zhang et al. (2023a); Huang et al. (2023). In addition, recent progress has seen the advent of individual-level agents that are utilized to simulate specific character behaviors, enhancing the realism and personalization of user-agent interactions Shao et al. (2023); Zhou et al. (2023); Wang et al. (2023d). This paper falls into simulation-oriented agents to build a human-like open-domain dialogue agent with memory retrieval and character analysis modules.
@@ 3 Method
In this section, we introduce the LD-Agent in detail with the framework shown in Figure 2. We first introduce the task definition of long-term dialogue in Section. 3.1. Consequently, we separately introduce the mechanism of event perception (Section. 3.2), dynamic personas extraction (Section. 3.3), and response generation (Section. 3.4).
@@ 3.2 Event Perception
The event memory module is designed to perceive historical events to generate coherent responses across intervals. In Figure 2, this event memory module is divided into two sub-modules that focus separately on long-term and short-term memory.
@@ 3.2.1 Long-term Memory
The long-term memory module aims to extract and encode events from past sessions. Specifically, this involves recording the occurrence times $t$ and brief summaries $o$ into representations that are stored in a low-cost memory bank $M_{L}=\{\phi(t_{j},o_{j})\mid j\in\{1,2,\dots,l\}\}$. Here, $\phi(\cdot)$ indicates the text encoder (e.g., MiniLM Wang et al. (2020)), and $l$ specifies the length of the memory bank. The encoded representations are then efficiently retrieved through an embedding-based mechanism, which enhances the accessibility of the stored memory.
@@
Different from previous agent approaches Park et al. (2023); Zhang et al. (2023a); Zhong et al. (2024) that entirely rely on LLM’s zero-shot ability to excavate and summarize events, we apply instruction tuning Wei et al. (2022a) to the event summary module, which can directly improve the event summary quality. Specifically, we rebuild the DialogSum dataset Chen et al. (2021), a large-scale dialogue summarization dataset, into the following format: (1) an introduction to the task background, (2) the related conversations that need to be understood, and (3) detailed summarization requests. These three parts serve as input prompts (see Appendix. D.1 for more details), combined with the original summaries from DialogSum as answers, and are jointly used to fine-tune the event summary module, thereby directly improving the quality of event summarization.
@@
To improve retrieval accuracy, we employ a retrieval mechanism that comprehensively considers semantic relevance, topic overlap, and time decay. Optimizing the retrieval accuracy of agent memory is challenging due to the difficulty in obtaining accurate memory retrieval data. Most existing methods Park et al. (2023); Zhang et al. (2023a) use event summaries as keys and context as queries, calculating the query-key semantic relevance score $s_{\text{sem}}$ to find relevant memories, which inevitably results in significant errors. To enhance retrieval reliability, we extract nouns from corresponding conversations with the summaries to construct a topic library $V$ and calculate topic overlap score $s_{\text{top}}$ by:
$s_{\text{top}}=\frac{1}{2}(\frac{\left|V_{q}\cap V_{k}\right|}{\left|V_{q}\right|}+\frac{\left|V_{q}\cap V_{k}\right|}{\left|V_{k}\right|}),$ (1)
where $V_{q}$, $V_{k}$ denote the topic noun set of query and key. Additionally, we refer to the recency coefficient used by Park et al. (2023) and apply an exponential decay function as time decay coefficient $\lambda_{t}=e^{{-t}/{\tau}}$ to reweight the overall retrieval score $s_{r}$, signified as Eq 2. $\tau$ is a temperature coefficient set to 1e+7 in our setting.
@@
To avoid retrieving inappropriate memory due to no suitable memories existing, we implement a semantic threshold $\gamma$ set to 0.5 in our setting. Only memories with semantic score $s_{\text{sem}}$ greater than $\gamma$ could be retrieved. If no appropriate memories are retrieved, “No relevant memory” will be returned. Eventually, the process of retrieving relevant memory can be denoted as $m=\psi(M_{L},\gamma)$.
@@ 3.2.2 Short-term Memory
The short-term memory module actively manages a dynamic dialogue cache $M_{S}=\{(t_{i},u_{i})|i=\{1,2,3,\dots,r_{c}\}\}$ with timestamps to preserve the detailed context of the current session. Upon receiving a new utterance $u^{\prime}$, the module first evaluates the time interval between the current time $t^{\prime}$ and the last recorded time $t_{r_{c}}$ in the cache. If this interval exceeds a threshold $\beta$ (600 seconds in our setting), the module triggers the long-term memory module to summarize the cached dialogue entries, creating new event records for storage in the long-term memory bank. Simultaneously, the short-term memory cache is cleared, and the new dialogue record $(t^{\prime},u^{\prime})$ is added to the cache. The mathematical expression of this process is given by:
@@
where $M^{\prime}_{L}$ denotes the updated long-term memory bank, $o=A(\cdot)$ is the event summary function, which process the accumulated dialogue in $M_{S}$.
@@ 3.3 Dynamic Personas Extraction
The persona module is pivotal in maintaining long-term persona consistency for both participants in a dialogue system. Drawing inspiration from prior work Xu et al. (2022b), we adopt a bidirectional user-agent modeling approach, utilizing a tunable persona extractor to manage long-term persona bank $P_{u}$ and $P_{a}$ for the user and agent, respectively. Specifically, we develop an open-domain, utterance-based persona extraction dataset derived from MSC Xu et al. (2022a). We enhance the persona extractor with LoRA-based instruction tuning, which allows for the dynamic extraction of personality traits during conversations. These traits are subsequently stored in the corresponding character’s persona bank. For utterances devoid of personality traits, the module outputs “No Trait”. Additionally, we employ a tuning-free strategy that harnesses the zero-shot capabilities of LLM models to directly extract personas based on prompts (see Appendix. D.2). To further improve the ability to excavate user personas without training, we adjust our reasoning strategy from direct reasoning to a Chain-of-Thought reasoning Wei et al. (2022b).
@@ 3.4 Response Generation
Upon receiving a new user utterance $u^{\prime}$, the agent integrates various inputs: retrieved relevant memories $m$, short-term context $M_{S}$, and the personas $P_{u}$ and $P_{a}$ for the user and agent, respectively. These combined inputs are fed into a response generator to deduce an appropriate response $r$, formulated as
@@
To enhance the agent’s ability for coherent and contextually appropriate responses, we develop a long-term, multi-session dialogue dataset, featuring dynamic retrieval memories, context, and personas sourced from the MSC and CC datasets for generator tuning. Specifically, for each sample, covering five sessions, we dynamically simulate the entire progression of the conversation. As each new utterance is introduced, the previously tuned modules for event summarization, persona extraction, and memory retrieval are utilized to collect the necessary context, retrieved memories, and both user and agent personas related to the utterance. This comprehensive data is then integrated into a response generation prompt (see Appendix. D.3). The original responses from the MSC and CC datasets are used as ground truth sentences.
@@ 4.1 Evaluation Settings
In this subsection, we briefly introduce the experimental dataset, evaluation metrics, and baseline models in our study. Detailed evaluation settings are elaborated in Appendix. A.
@@
ChatGLMLDA 7.42 2.46 20.04 7.47 2.40 19.50 7.52 2.32 19.55 7.36 2.37 19.16
Tuning $\text{ChatGLM}_{\text{LDA}}^{*}$ 10.70 5.63 23.31 10.03 5.12 21.55 9.07 4.06 20.19 8.96 4.01 19.94
@@
Table 1: Experimental results of the automatic evaluation for response generation on MSC and CC. * denotes using annotations as personas.
Datasets.
To investigate the effectiveness of LD-Agent on long-term dialogue scenarios, extensive experiments are conducted on two illustrative multi-session datasets, MSC Xu et al. (2022a) and CC Jang et al. (2023), each comprising 5 sessions with approximately 50 conversational turns per sample. The experiments cover model independence assessment, module ablation, persona extractor analysis, and cross-domain evaluation. Additionally, to evaluate the transferability of the LD-Agent, we apply our method to the Ubuntu IRC benchmark Hu et al. (2019), a dataset known for its multiparty interaction tasks.
Metrics.
Our evaluation combines both automatic and human assessments to thoroughly investigate the effectiveness of LD-Agent. For automatic evaluation, we use three widely used standard metrics: BLEU-N (BL-N) Papineni et al. (2002), ROUGE-L (R-L) Lin (2004), and METEOR (MET) Banerjee and Lavie (2005) to measure the quality of response generation. Additionally, accuracy (ACC) is employed to evaluate the classification performance of the persona extractor. In human evaluation, we measure topic coherence across sessions, interaction fluency, and user engagement using the metrics of coherence, fluency, and engagingness, respectively.
Baselines.
To demonstrate the effectiveness and model independence of LD-Agent, we deploy it on multiple platforms and models. Specifically, the LLM-based models (online model: ChatGPT; offline model: ChatGLM Zeng et al. (2023)) and traditional language models (BlenderBot Roller et al. (2021), and BART Lewis et al. (2020)) are employed as our baselines. In our experiments, The notation “ModelLDA” denotes models that incorporate the LD-Agent framework, while “Model” refers to the original baseline models without LD-Agent. Additionally, we also utilize HAHT Zhang et al. (2022), the previous state-of-the-art model in long-term dialogue task, as a contrast.
4.2 Evaluation Pipeline
Automatic Evaluation Pipeline.
For the automatic evaluation, We first utilize the first session in MSC or CC to initialize the conversation. Afterward, we calculate generation accuracy for each utterance. For instance, in a 30-turn dialogue where 15 utterances come from Speaker B, who will be role-played by the Agent. Accuracy is calculated based on all of these 15 utterances. To simulate a realistic dialogue scenario, before generating each utterance, we first input all the preceding ground-truth conversations into the LD-Agent framework to simulate the real historical interaction process. During the simulation, personas and memories are automatically extracted to assist in generating the current utterance. Additionally, since the MSC dataset has annotated personas, we also evaluated using these annotations as personas instead of extracting them, marked with *.
Human Evaluation Pipeline.
In Section 4.6, we conduct human evaluation on memory retrieval and response generation. Specifically, we employ 8 students to assist with the assessment and evaluate LD-Agent on two tasks: memory retrieval evaluation and response generation evaluation. The detailed guidelines of human evaluation are shown in Figure 4 of Appendix.